Write a recursive formula for the sequence 15,26,48,92,180,.......Then find the next term.

Answer by stanbon(75887)

15,26,48,92,180,.......
---
15
15+1*11 = 26
26 + 2*11 = 48
48 + 4*11 = 92
92 + 8*11 = 180
------------------------
180 + 16*11 = 356
========================
Cheers,
Stan H.